a carpenter bought 10 item. some cost 40 and the remaining 30.50. how many of each were bought if the total was 362

Answer by stanbon(75887)

a carpenter bought 10 item. some cost 40 and the remaining 30.50. how many of each were bought if the total was 362
-----
quantity equation: f + t = 10
value equation:: 40f +30.5t = 362
---------------------------------------
Modify the equations:
400f + 400t = 4000
400f + 305t = 3620
--------------------------
Subtract and solve for "t":
95t = 380
---
t = 4 (# of 30.5 items)
---
Solve for "f":
f + t = 10
f = 4 = 10
f = 6 (# of 40 items)
========================
Cheers,
Stan H.